A room contains 96 chairs. The number of chairs per row is 4 more than the number of rows. Find the number of rows and the number of chairs per row.

C+R = 96 Chairs
4c + R= 96

can you please help me with this problem..THanks

Found 2 solutions by ankor@dixie-net.com, checkley77:
Answer by ankor@dixie-net.com(22740)

A room contains 96 chairs.
The number of chairs per row is 4 more than the number of rows.
Find the number of rows and the number of chairs per row
:
Let c = no. of chairs per row
Let r = no. of rows
:
r * c = 96
:
"The number of chairs per row is 4 more than the number of rows.
c = r + 4
:
Replace c with (r+4) in the 1st equation
r(r+4) = 96
:
r^2 + 4r - 96 = 0; a quadratic equation
Factors to
(r+12)(r-8) = 0
The positive solution
r = 8 rows
and obviously
12 chairs in each row

Answer by checkley77(12844)

r*c=96
c=r+4
r(r+4)=96
r^2+4r-96=0
(r+12)(r-8)=0
r-8=0
r=8 number of rows.
c=8+4=12 chairs per row.
Prof:
8*12=96
96=96